# Over-the-Air Federated Adaptive Data Analysis: Preserving Accuracy via Opportunistic Differential Privacy

### v1

@@ IV Point-to-Point ADA in the Presence of a Gaussian Channel
At the first step, to make our idea clear, we start with the case in which there is only one EP, answering the PS’s adaptive queries via a Gaussian channel. More exact, at the $i$’th question-answering round of this interactive procedure, PS sends $q_{i}(\cdot)$ through an error-free channel and receives $a_{i}$ as the answer from EP through a Gaussian channel with noise variance of $\sigma^{2}$. Conventionally, noise is considered as a limiting factor in the communication procedures. But in this work we opportunistically employ it as a constructive element. As noted in I-A we aim to exploit the intrinsic channel noise as the randomization noise samples in the Gaussian answering mechanism. In this scheme, at each round, such as round $i$, the EP computes the empirical average of the query as
@@
then multiplies it by $A_{\mathsf{t}}$. The resulting signal is broadcast without additional coding, using analog Pulse Amplitude Modulation (PAM) with a maximum amplitude of $A_{\mathsf{t}}$. Using a matched filter, PS detects a noisy version of the transmitted signal. Without any hard decision, he considers the detected signal as $a_{i}$, i.e. the answer of his asked query. So we have:
@@
where $z_{i}\sim\mathcal{N}(0,\sigma^{2})$ is the sampled Gaussian channel noise at the receiver. Assuming that the channel condition remains unchanged during the whole ADA procedure, it’s clear that $n_{i}$’s are independent and identically distributed. So, the proposed scheme completely resembles the Gaussian answering mechanism in ADA and we can apply theorem III.1 to this case. Actually we use this theorem to set the tunable communication parameters in a way that the desired accuracy level is met. We assume that $(\alpha,\beta)$ has been given and the dataset size, $n$, is fixed. Based on this parameters and as a function of $A_{\mathsf{t}}$, we can derive the maximum number of allowable queries, for which still (2) remains valid.
In the basic model of ADA, it’s assumed that the queries maximum value is fixed and equal to 1. There, the variance of the artificial randomization noise is calibrated in a way that the best performance is achieved. On the other hand, in our proposed model, the noise variance is fixed and determined by the physical later conditions. Hence, here we instead tune the queries maximum amplitude, i.e. $A_{\mathsf{t}}$, for the best performance achievement. Also note that $\alpha$ in (2) is considered under the assumption that $|q_{i}(.)|\leq 1$. Nonetheless, we keep using (2) by normalizing the answers maximum values to 1, when investigating this relation. More exact, the validity condition of (2) for the situation $(\alpha=\alpha_{0}\cdot A_{0},\beta=\beta_{0},A_{\mathsf{t}}=A_{0},\sigma=\sigma_{0})$ is completely equivalent with the situation $(\alpha=\alpha_{0},\beta=\beta_{0},A_{\mathsf{t}}=1,\sigma=\sigma_{0}/A_{0})$. So, by tuning $A_{\mathsf{t}}$ we implicitly calibrate the noise variance in the equivalent normalized model.
@@
Also, in figure 4 the maximum achievable number of queries is drawn versus $n$. In this figure, for each $n$, the best $\sigma/A_{\mathsf{t}}$ that maximizes $k$ is considered. As depicted in the figure, $k_{m}ax$ follows a quadratic growth in $n$. This observation confirms the well known fact in the ADA problem that for the worst case setup, till $O(n^{2})$ number of queries can be $(\alpha,\beta)$-accurately be answered [20].
@@
Furthermore, figure 3 shows that as $n$ increases, the optimum value of $\sigma/A_{\mathsf{t}}$ very slowly shifts to the left. This optimum point is the intersection of the quadratically increasing left curve and the exponentially decreasing right curve, , see $k_{1}$ and $k_{2}$ in (8). For our considered dataset sizes, optimum $\sigma/A_{\mathsf{t}}$ approximately lies in $[0.008,0.01]$. Furthermore, when $\sigma/A_{\mathsf{t}}$ tends to zero, the number of accurate queries reduce to values close to zero, too. In this low-noise regime the amount of randomization is not sufficient for ensuring that the dataset will not leak. Consequently, in this case by asking only few queries the analyst can overfit to the dataset and violate the accuracy criterion (2). Hence, in contrast to ordinary communication scenarios, increasing the power transmission may cause non-efficiency in this setup. On the other hand, computation shows that in order to have a non-zero number of $(\alpha,\beta)$-accurate query-answers by this scheme, $\sigma/A_{\mathsf{t}}$ should be lower than 0.017 or equivalently $A_{\mathsf{t}}$ should be approximately 60 times greater than the standard deviation of the noise. This ratio for the optimum situation lies in $[100,125]$. Equivalently, a minimum SNR of 32 dB is required and optimum SNR is in the range [37, 38.9] dB. However, in high-noise regime, considering practical limitations on the transmission power, improving $k$ may not be straight-forward. We show in the next section that federation of some edge answering points can remedy this situation.
@@ V Distributed ADA in the Presence of a Gaussian Channel
In the previous section we considered the point-to-point scenario to make the idea of usefully exploiting the channel noise clear. We showed that by calibrating the amplitude of transmitted signal, achievable number of accurately answered queries can be optimized. Furthermore, we see that optimum $A_{\mathsf{t}}$ depends on the dataset size, $n$. On the other hand, although we’ve succeeded theoretically to utilize the noise as a part of adaptive query answering procedure, but the required SNR range may not be manageable in some of communication situations. This concern becomes more crucial when the answerer is a personal device or a low-power sensor at the edge of network with a long distance from the receiver. As we mentioned previously, this can be a reasonable use case of our proposed setup. In addition, we observed that for a fixed $A_{\mathsf{t}}$ the required dataset size, i.e. $n$, grows by $O(\sqrt{k})$. So, to guarantee more accurate (or private) rounds of data analysis, the edge point needs to answer the queries based on larger datasets. However, it may be impractical for him to gather or store large samples. In this section we show that both of the power and data limitations can be handled in the distributed scenario.
We consider a situation that $L$ Edge Points simultaneously transmit their answers to adaptive queries of the Central point through an analog Gaussian multiple access channel (MAC). For simplicity, let’s assume that the dataset sizes of EPs are the same, i.e. $n_{l}=n_{0}$ for any $l\in\{1,2,\dots,L\}$. Similar to the previous section, each EP, like $\mathcal{M}_{l}$, computes $q_{i}(S_{l})=\frac{1}{n_{0}}\sum_{j=1}^{n_{0}}{q_{i}(x_{j,l})}$ and transmit $A_{\mathsf{t}}q_{i}(S_{l})$ without any coding. Because of the additive nature of the air interface, the receiver gets a noisy version of the aggregated signals. So, the received signal at the CP will be $a_{i}=A_{\mathsf{t}}\sum_{l=1}^{L}{q_{i}(S_{l})}+z_{i}$, in which $z_{i}\sim\mathcal{N}(0,\sigma^{2})$. In order to make an equivalence with the point-to-point case, we consider $a_{i}/L$ in the following. By substituting $q_{i}(S_{l})$, we will have:
@@
The above equation completely resembles a point-to-point scenario with $n_{\text{eq}}=Ln_{0}$, $\sigma_{\text{eq}}=\frac{\sigma}{L}$ and $A_{T,eq}=A_{\mathsf{t}}$. We see that by this cooperative transmission scheme, both the equivalent dataset and the noise variance is improved by a factor of $L$. By considering these equivalent parameters, the previous section results can be applied in this scenario, too. Therefore, we can use (7) with the aforesaid $n_{\text{eq}}$ and $\sigma_{\text{eq}}$. In this case, we fix $\sigma/A_{\mathsf{t}}=0.5$ and derive $k$ as a function of $L$, for different values of $n$. Considering the previous section, we know that $L$ should be large enough to make $\sigma_{\text{eq}}$ less than $0.017$ and consequently, attain a positive value for $k$. Figure 5 shows the results for this scenario. The parameters $\alpha$ and $\beta$ are the same with the previous scenario.
@@
We see that for each $n$, before a critical number of EPs, $k$ grows exponentially in $L$, which is due to under-leakage regime and is related to $k_{2}$ in (8). This critical $L$ shifts very slowly in $n$ to the right. For each $n$, after this $L$, growth of $k$ becomes very slow such that in figure 5 it seems constant. The reason is that after the critical $L$, $k$ obeys from $k_{1}$ function in (8). As we noted previously, $k_{1}$ can be approximated by $\hat{k_{1}}$ in (10). In this function, the quadratic term is strongly dominant for large $n$. In this quadratic term, $n^{2}\sigma^{2}$ is determining. On the other hand, in the distributed scenario we have
$n_{\text{eq}}\sigma_{\text{eq}}=(Ln)(\sigma/L)=\sigma,$
@@
V-A Improving number of queries in EP counts
Considering our descriptions on the figure 3 about the optimum value of $\sigma/A_{\mathsf{t}}$, EPs in the distributed scenario need to decrease $A_{\mathsf{t}}$ with respect to $L$, after that $L$ passes the critical value. For large $n_{\text{eq}}$, optimum $\sigma_{\text{eq}}/A_{\mathsf{t}}$ is about $0.008$. The exact value can be derived by setting $k_{1}=k_{2}$ in (7). Naming this value by $s_{\text{opt}}(n_{\text{eq}})$, optimum $\sigma/A_{\mathsf{t}}$ will be $Ls_{\text{opt}}(n_{\text{eq}})$. In other words, EPs should tune their $A_{\mathsf{t}}$ equal to $\frac{\sigma}{Ls_{\text{opt}}(Ln)}$. As we noted, in the previous section, $s_{\text{opt}}$ changes very slowly. Hence, we can approximately say that in optimum regime $A_{\mathsf{t}}\propto 1/L$. Figure 6 depicts the resulted improvement in the growth of $k$, after optimizing $A_{\mathsf{t}}$.
@@ VI Future Work and Conclusion
In this work, after making a previous result in ADA more proper for our analyses, we showed that Gaussian channel noise can be used as an opportunity when the basic model of ADA is considered. Of course, in this situation minimum required SNR was considerable. Then, we showed that the federation of the edge answering points can improve the equivalent noise variance and total dataset size by a linear factor when the shared medium is a Gaussian MAC.
One important extension for further study can be considering the fading channel, which is a more realistic model. In this situation, each EP’s transmission experiences a multiplicative distortion before receiving by the CP. In the other word, the received signal during the i’th answering round will be as follows
@@
in which $X_{il}$ is the transmitted signal by the $l$’th EP and $h_{il}$ is the fading coefficient of the wireless channel from this EP to the CP. Based on various scenarios of Channel State Information (CSI) knowledge, different scenarios can be studied: CSI known only at the transmitter, CSI known only at the receiver, and CSI not known at either. Overally, a main challenge in the fading scenario is that the unweighted aggregation of transmissions is not intrinsically prepared anymore.

### v2

@@ IV Point-to-Point ADA in the Presence of a Gaussian Channel
To clarify our idea, we begin with the case where there is only one EP answering the PS’s adaptive queries via a Gaussian channel. Specifically, at the $i$’th question-answering round of this interactive procedure, the PS sends $q_{i}(\cdot)$ through an error-free channel and receives $a_{i}$ as the answer from the EP through a Gaussian channel with noise variance $\sigma^{2}$. As noted in I-A, we aim to exploit the intrinsic channel noise as the randomization noise samples in the Gaussian answering mechanism. In this scheme, at each round, such as round $i$, the EP computes the empirical average of the query as
@@
and then multiplies it by $A_{\mathsf{t}}$. The resulting signal is broadcast without additional coding, using analog Pulse Amplitude Modulation (PAM) with a maximum amplitude of $A_{\mathsf{t}}$. Using a matched filter, the PS detects a noisy version of the transmitted signal. Without applying any hard decision, it interprets the detected signal as $a_{i}$, the answer to its query. Therefore, we have:
@@
where $z_{i}\sim\mathcal{N}(0,\sigma^{2})$ is the sampled Gaussian channel noise at the receiver. Assuming that the channel condition remains unchanged during the whole ADA procedure, it is clear that $n_{i}$’s are independent and identically distributed. Thus, the proposed scheme completely resembles the Gaussian answering mechanism in ADA, allowing us to apply Theorem III.1 to this scenario. We use this theorem to adjust the tunable communication parameters such that the desired accuracy level is achieved. We assume that $(\alpha,\beta)$ has been given, and the dataset size, $n$, is fixed. Based on these parameters and as a function of $A_{\mathsf{t}}$, we can derive the maximum number of allowable queries for which (2) remains valid.
In the basic ADA model, the maximum value of the queries is assumed to be fixed at 1, and the variance of the artificial randomization noise is adjusted to achieve optimal performance. In contrast, in our proposed model, the noise variance is dictated by the physical layer conditions and is beyond our control. Instead, we tune the maximum amplitude of the query, denoted as $A_{\mathsf{t}}$, to achieve the best performance. Also, note that $\alpha$ in (2) is considered under the assumption that $|q_{i}(\cdot)|\leq 1$. Nonetheless, we continue using (2) by normalizing the maximum values of the answers to 1 when investigating this relation. More specifically, the validity condition of (2) for the situation $(\alpha=\alpha_{0}\cdot A_{0},\beta=\beta_{0},A_{\mathsf{t}}=A_{0},\sigma=\sigma_{0})$ is entirely equivalent to the situation $(\alpha=\alpha_{0},\beta=\beta_{0},A_{\mathsf{t}}=1,\sigma=\sigma_{0}/A_{0})$. Thus, by tuning $A_{\mathsf{t}}$, we implicitly calibrate the noise variance in the equivalent normalized model.
@@
Additionally, in Figure 4, the maximum achievable number of queries is plotted against $n$. For each value of $n$, the optimal $\sigma/A_{\mathsf{t}}$ that maximizes $k$ is selected. As shown in the figure, $k_{\text{max}}$ exhibits quadratic growth with respect to $n$. This observation aligns with the well-known result in the ADA literature that, in the worst-case setup, up to $\mathcal{O}(n^{2})$ queries can be answered $(\alpha,\beta)$-accurately [20].
@@
Furthermore, Figure 3 shows that as $n$ increases, the optimum value of $\sigma/A_{\mathsf{t}}$ very slowly shifts to the left. This optimal point corresponds to the intersection of the quadratically increasing left curve and the exponentially decreasing right curve, as illustrated by $k_{1}$ and $k_{2}$ in (8). For the dataset sizes considered in this study, the optimal value of $\sigma/A_{\mathsf{t}}$ approximately falls within the range $[0.008,0.01]$. Furthermore, as $\sigma/A_{\mathsf{t}}$ approaches zero, the number of accurate queries also declines to near-zero levels. In this low-noise regime, the level of randomization becomes insufficient to prevent data leakage. As a result, the analyst can overfit to the dataset and violate the accuracy criterion in (2) with only a few queries. Thus, unlike standard communication scenarios, increasing transmission power in this setup can lead to inefficiency. On the other hand, computation shows that in order to have a non-zero number of $(\alpha,\beta)$-accurate query-answers by this scheme, $\sigma/A_{\mathsf{t}}$ should be lower than 0.017 or equivalently $A_{\mathsf{t}}$ should be approximately 60 times greater than the standard deviation of the noise. This ratio, in the optimum situation, lies within the range of $[100,125]$. Equivalently, a minimum SNR of 32 dB is required, with the optimum SNR falling between 37 and 38.9 dB. However, in the high-noise regime, given practical limitations on transmission power, improving $k$ may not be straightforward. In the next section, we demonstrate how a federation of edge answering points can address this issue."
@@ V Distributed ADA in the Presence of a Gaussian Channel
In the previous section, we examined the point-to-point scenario to clarify the concept of effectively exploiting the channel noise. We showed that calibrating the amplitude of the transmitted signal can optimize the achievable number of accurately answered queries. Furthermore, we see that optimum $A_{\mathsf{t}}$ depends on the dataset size, $n$.
On the other hand, although we have theoretically succeeded in utilizing noise as part of the adaptive query-answering procedure, the required SNR range may not always be feasible in certain communication scenarios. This challenge is particularly significant when the answerer is a personal device or a low-power sensor located at the network’s edge and operating at a considerable distance from the receiver. As previously mentioned, this represents a plausible use case for our proposed setup. Additionally, we observed that for a fixed $A_{\mathsf{t}}$, the required dataset size $n$ grows at a rate of $\mathcal{O}(\sqrt{k})$. Consequently, ensuring more accurate (or private) rounds of data analysis would require the EP to answer queries based on increasingly larger datasets. However, it may be impractical for the EP to gather or store such large datasets. In this section, we demonstrate that both power and data limitations can be addressed within a distributed framework.
We consider a situation where $L$ EPs simultaneously transmit their answers to the CP’s adaptive queries through an analog Gaussian multiple access channel (MAC). For simplicity, we assume that the dataset sizes of all EPs are identical, i.e., $n_{l}=n_{0}$ for all $l\in{1,2,\dots,L}$. Similar to the previous section, each EP, like $\mathcal{M}_{l}$, computes $q_{i}(S_{l})=\frac{1}{n_{0}}\sum_{j=1}^{n_{0}}{q_{i}(x_{j,l})}$ and transmit $A_{\mathsf{t}}q_{i}(S_{l})$ without any coding. Because of the additive nature of the air interface, the receiver gets a noisy version of the aggregated signals. So, the received signal at the CP will be $a_{i}=A_{\mathsf{t}}\sum_{l=1}^{L}{q_{i}(S_{l})}+z_{i}$, where $z_{i}\sim\mathcal{N}(0,\sigma^{2})$. To establish equivalence with the point-to-point case, we consider $a_{i}/L$ in the following. Substituting $q_{i}(S_{l})$ yields:
@@
The above equation completely resembles a point-to-point scenario with $n_{\text{eq}}=Ln_{0}$, $\sigma_{\text{eq}}=\frac{\sigma}{L}$ and $A_{T,eq}=A_{\mathsf{t}}$. We observe that with this cooperative transmission scheme, both the equivalent dataset size and the noise variance are improved by a factor of $L$. Using these equivalent parameters, the results from the previous section can be applied to this scenario as well. Thus, we can use (7) with the corresponding $n_{\text{eq}}$ and $\sigma_{\text{eq}}$. In this case, we fix $\sigma/A_{\mathsf{t}}=0.5$ and derive $k$ as a function of $L$ for different values of $n$. Based on the results from the previous section, we know that $L$ must be sufficiently large to ensure that $\sigma_{\text{eq}}$ is less than $0.017$, thereby yielding a positive value for $k$. Figure 5 illustrates the results for this scenario. The parameters $\alpha$ and $\beta$ remain the same as in the previous case.
@@
We observe that for each $n$, before reaching a critical number of EPs, $k$ grows exponentially with $L$. This is due to the under-leakage regime and is related to $k_{2}$ in (8). The critical value of $L$ shifts slowly to the right as $n$ increases. After this critical $L$, the growth of $k$ becomes very slow, and in Figure 5, it appears nearly constant. This is because, beyond the critical $L$, $k$ follows the $k_{1}$ function from (8). As mentioned earlier, $k_{1}$ can be approximated by $\hat{k_{1}}$ in (10). In this function, the quadratic term becomes strongly dominant for large $n$. In this quadratic term, $n^{2}\sigma^{2}$ is determining. On the other hand, in the distributed scenario, we havebb
$n_{\text{eq}}\sigma_{\text{eq}}=(Ln)(\sigma/L)=\sigma,$ (14)
@@
V-A Improving $k$ in $L$
Considering the descriptions in Figure 3 about the optimum value of $\sigma/A_{\mathsf{t}}$, EPs in the distributed scenario need to decrease $A_{\mathsf{t}}$ as $L$ exceeds the critical value. For large $n_{\text{eq}}$, the optimum value of $\sigma_{\text{eq}}/A_{\mathsf{t}}$ is approximately $0.008$. The exact value can be derived by setting $k_{1}=k_{2}$ in (7). Denoting this value as $s_{\text{opt}}(n_{\text{eq}})$, the optimum value of $\sigma/A_{\mathsf{t}}$ will be $Ls_{\text{opt}}(n_{\text{eq}})$. In other words, EPs should tune their $A_{\mathsf{t}}$ to $\frac{\sigma}{Ls_{\text{opt}}(Ln)}$. As noted in the previous section, $s_{\text{opt}}$ changes very slowly. Therefore, we can approximately state that in the optimum regime, $A_{\mathsf{t}}\propto 1/L$. Figure 6 illustrates the resulting improvement in the growth of $k$ after optimizing $A_{\mathsf{t}}$.
@@ VI Future Work and Conclusion
In this work, after refining a previous result in ADA to better suit our analyses, we showed that Gaussian channel noise can be leveraged as an opportunity when the basic model of ADA is considered. Of course, in this situation, the minimum required SNR was significant. Then, we demonstrated that the federation of edge answering points can linearly improve the equivalent noise variance and total dataset size when the shared medium is a Gaussian MAC.
One important extension for further study could involve considering the fading channel, which represents a more realistic model. In this case, each EP’s transmission undergoes a multiplicative distortion before being received by the CP. In other words, the received signal during the $i$-th answering round would be as follows:
@@
in which $X_{il}$ represents the signal transmitted by the $l$-th EP, and $h_{il}$ denotes the fading coefficient of the wireless channel between this EP and the CP. Based on the availability of CSI, various scenarios can be analyzed: CSI known only at the transmitter, CSI known only at the receiver, and CSI unknown at both ends. A primary challenge in the fading scenario is that the unweighted aggregation of transmissions is no longer inherently feasible.